OpenStudy (amtran_bus):

If M(5, 4) is the midpoint of the line segment AB, and if A has coordinates (−5, −1), find the coordinates of B.

OpenStudy (amtran_bus):

Is it (15,9)?

OpenStudy (anonymous):

the formula for determining the coordinates of a midpoint is x = (xA + xB)/2 y = yA + yB)/2 you are right!

OpenStudy (amtran_bus):

Really? WOW! I usually don't get it the first time.

OpenStudy (anonymous):

Be confident

OpenStudy (amtran_bus):

Good advice.

OpenStudy (amtran_bus):

Thank you.

OpenStudy (anonymous):

My pleasure